virtual method; a=10t b=5m
Soru 5 (25 p) Şekilde verilen Virtüel iş Yönte EI b ЗЕТ 2EI КІ 72 2b
Answer #1

Humber of Reclandants: 7: 4-3=1 Taking horizontal peachon at L as unkown Primary analysis: (M) lot EFX=0 som ad Kx=104 sm EMLCompaitibích sidity equations 2638.89 1034.72 At fix L2=0 =) · Læs =20557 10 faco ka=10-2055= foust EML=0 10X550 Kyx kyrio ky